Oxygen Supply in Submarines Nuclear submarines can stay under water nearly indefinitely because they can produce their own oxygen by the electrolysis of water.

How many liters of O2 at 298 K and 1.00 bar are produced in 1.50 hr in an electrolytic cell operating at a current of 0.0200 A?
Answer #1

Electrolysis of water is given below:

At cathode: 4H+(aq) + 4e- -----------> 2H2(g)

At Anode: 2H2O(l) -------> O2(g) + 4H+ + 4e-

1.5 hrs = 5400 sec

Total charge = Q = current * time (in sec)

Q = 0.02*5400 = 108 coulombs

We know that,

96500 Coulombs charge is carried by 1 mole electrons

So, 108 Coulomb charge is carried by 102/96500 moles of electrons.

So, moles of e- = 108/96500 = 0.00112 moles

From reaction at Anode, it is clear that 4 moles of e- produces 1 mole of O2

So, moles of O2 (n)= 0.00112/4 = 0.00028 moles

P = 1 bar = 0.987 atm , T = 298K , R = 0.0821 L atm/mol K

Let volume of O2 is V liters

Also,

PV = nRT then V = nRT/P

V = 0.00028*0.0821*298/0.987 = 0.00695 liters

OR

V = 0.007 Liters ...... Answer